OpenStudy (anonymous):

solve this equation. (-9)(-9)4 = m m=

OpenStudy (anonymous):

\(\Large m=324\)

OpenStudy (anonymous):

\[m = (-9)(-9)4\]This is the same as \[m = 9*9*4\]because of the double negatives that becomes positive. If you ever are in doubt, you can use the calculator to confirm your answer. 9*9 = 81. 81*4 = 324. Remember that knowing "how" is more important than "what". :)